Shift Summary for DAY shift, 11/19/99

Yee Bob Hsiung (hsiung@fnal.gov)
Fri, 19 Nov 1999 15:32:56 -0600

Shift Summary for DAY shift, 11/19/99

On Shift: Viv O'Dell, Bob Hsiung

Hours of beam available at all intensities: 8.4
Hours of beam available at nominal intensity: 8.3
Hours of E799 data (special runs not included): 7.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: Run changes, Run abort, drt_fresh_starts,
bad tape/drive?, etc. A few missing spills from TeVatron.

========================================================================
Runs ended this shift:

Run 15007 - 22.5M events, 376K 3pi0D.
Run 15008 - 11 spills, DAQ hung, overwritten.
Run 15011 - ~5 spills, abort by itself, overwritten.
Run 15013 - in progress. (swapped a new tape for UPB159)

Tapes brought to FCC this shift:
(including tapes for online split)

UPA-C148-157 and UPS050-52.

========================================================================
Accesses during shift:

none

========================================================================
Overview of shift:
E799 datataking until 10am Tapes full (though only 80%),
Update CSQ615007. New set of tapes.
Then had trouble with new run, DAQ hung after 11 spills, multiple
drt_fresh_starts etc. and the next run aborted by itself
after about 5 spills. Suspecting a bad tape/drive on ktev4 drv2.
Replaced a new tape, then followed by smooth sailing.

SEM varied between 8.5e12 and 8e12, slowly dropping over the day shift.

Labeled some blank tapes for data taking.
Viv madly pounding on kpasa/ksera to get more space for user area.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15007 10:00 "Wow, I am full." Run stopped. (tapes are 80% full)
Paged Ed for updating CsI constants. Change tapes.
Init new run.
15008 1 10:15 Start run.
11 DAQDEAD_SEC=39.3, livetime is 0. ktev4 hung.
10:35 Stop run, but run won't stop.
drt_fresh_start and overwrite this run.
after several tries and mutiple logout and login.
15011 1 11:05 Finally succeed starting a new run.
11:17 "Shihi, abort abort abort" run aborted by Taku.
ktev4 drv2 (UPB159) light still blinking, suspect
a bad tape/drive, swap a new tape.
Overwrites this run again.
15013 1 11:45 Run started after drt_fresh_start.

Smooth run since and only had two kfrends
(one requested by Valeri, the other due to
trigger stopped)
13:15 Call MCR for more intensity, since it was dropping
slowly over the day.
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift